OpenStudy (anonymous):

Find the sum of the first eight terms: -800 + (-200) + (-50) + (-12.5) + … (Points : 1) -1062.5 -1066.60 -1066.68 -1066.65

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Can you tell me the relationship between the terms? In other words, how do we find those missing terms to add them up?

OpenStudy (whpalmer4):

is this an arithmetic or geometric sequence?

OpenStudy (anonymous):

geometric

OpenStudy (whpalmer4):

Okay, and the constant ratio?

OpenStudy (anonymous):

Ummm I don't know

OpenStudy (whpalmer4):

well, if the first term is -800, and the second term is -200, what number do we multiply with -800 to get -200?

OpenStudy (anonymous):

4?

OpenStudy (whpalmer4):

That's the common ratio or constant ratio or geometric ratio

OpenStudy (anonymous):

Oh okay :)

OpenStudy (whpalmer4):

-800 * 4 = -200?

OpenStudy (anonymous):

The answer is -1066.65

OpenStudy (whpalmer4):

well, that's nice, but do you want to know how to get it, or not? again, I ask for confirmation that you think that -800 * 4 = -200

OpenStudy (whpalmer4):

Personally, I don't think -800 * 4 = -200. I think -800 * 1/4 = -200...

OpenStudy (anonymous):

Actually that is correct. I divided by 4 to get my answer which is the same as multiplying by 1/4